# On the Kantorovich contraction of Markov semigroups

### v1

@@ 3.4.4 Example 2: A fixed point theorem
Let $F:S\rightarrow S$ be a mapping such that (3.26) holds for some $x_{0}\in S$ and parameters $\lambda(x_{0})\in]0,1[$ and $c(x_{0})\geq 0$. Next, we let $\delta>0$, $W(x)=\frac{1}{2}\exp{(\delta\|x\|)}$ and assume that there is an $r_{0}>0$ such that, for all $r\geq r_{0}$, there exists an $\alpha(r)\in]0,1[$ such that
@@ 3.5.1 A general diffusion model
We let $S=\mathbb{R}^{n}$, $\psi_{S}(x,y)=\|x-y\|$ and we consider below the continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ of a diffusion flow $y\in S\mapsto X_{t}(y)$ defined by
@@
In the above display, $B_{t}$ is a $n_{1}$-dimensional Brownian motion starting at the origin for some $n\geq 1$ and $b$ is a differentiable drift function from $\mathbb{R}^{n}$ into itself satisfying, for some constants $a>0$ and $c\geq 0$,
@@
Moreover, $\tau$ stands for some $(n\times n_{1})$-matrix valued continuous function such that, for all $x\in S$, the matrix $\tau(x)\tau(x)^{\top}$ is invertible and the map $x\mapsto\|\tau(x)\|/\|x\|$ vanishes at infinity. In what follows, for any $x\in S$ we let
$\|\tau(x)\|_{F}^{2}:=\mbox{\rm Tr}(\tau(x)\tau(x)^{\top})\leq n_{1}\|\tau(x)\|_{2}^{2}\quad\text{ with }\|\tau(x)\|_{2}^{2}=\lambda_{\max}(\tau(x)\tau(x)^{\top})$
@@ 3.5.2 The overdamped Langevin diffusion model
An important class of diffusion processes arising in statistical physics and applied probability is the overdamped Langevin diffusion flow $X_{t}(y)$, defined as in (3.30) with $(n_{1},\tau(x))=(n,\sigma I)$ for some $\sigma>0$ and the drift function $b(x)=-\gamma\nabla U(x)$ associated with some parameter $\gamma>0$ and smooth potential function $U$ satisfying the following condition
@@
$W(x):=(1/2)\exp(\delta\|x\|)\Longrightarrow\sup_{0\leq h^{\prime}\leq 1}\|\widetilde{Q}_{h^{\prime}}(W)/W\|<\infty\text{ and }\widetilde{Q}_{h}(W)/W\leq c_{h}/W^{\delta_{h}}.$ (3.34)
@@ 4 Proofs
showing that it $D_{\psi}(\mu_{1},\mu_{2})>0$ then
@@
Let $x,y\in S$ be such that $x\neq y$ and $r>r_{\epsilon}\vee r_{0}$ be fixed in what follows. Then, under the assumption that (H2) holds with the semi-distance $\kappa=\varphi_{0}$, we have
$\|\delta_{x}P-\delta_{y}P\|_{\tiny tv}\leq 1-\alpha(r)<1.$ (4.12)
In what follows the functions $(\varpi_{\rho},\varphi_{\rho})$ indexed by $\rho>0$ are as defined in (2.3) and (2.4).
@@
$\begin{split}\ell_{\varphi_{\rho}}(P)(x,y)&\leq\ell_{\varpi_{\rho}}(P)(x,y)=\frac{1+\rho P(V)(x)+\rho P(V)(y)}{1+\rho\varpi_{V}(x,y)}\\ &\leq\frac{1+\rho\big(1+\epsilon\varpi_{V}(x,y)\big)}{1+\rho\varpi_{V}(x,y)}\\ &=\epsilon+\frac{1+\rho-\epsilon}{1+\rho r}=1-(1-\epsilon)~\frac{\rho r}{1+\rho r}~\left(1-\frac{r_{\epsilon}}{r}\right)\end{split}$ (4.13)
where the first inequality uses the assumption that (H1) holds with $c=1/2$ and the last inequality uses (4.12). Simple calculations show that

### v2

@@ 3.4.4 Example 2: A fixed point theorem
Let $F:S\rightarrow S$ be a mapping such that (3.26) holds for some $x_{0}\in S$ and parameters $\lambda(x_{0})\in]0,1[$ and $c(x_{0})\geq 0$. Next, we let $\delta>0$, $W(x)=\exp{(\delta\|x\|)}$ and assume that there is an $r_{0}>0$ such that, for all $r\geq r_{0}$, there exists an $\alpha(r)\in]0,1[$ such that
@@ 3.5.1 A general diffusion model
We let $S=\mathbb{R}^{m}$, $\psi_{S}(x,y)=\|x-y\|$ and we consider below the continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ of a diffusion flow $y\in S\mapsto X_{t}(y)$ defined by
@@
In the above display, $B_{t}$ is an $m_{1}$-dimensional Brownian motion starting at the origin and $b$ is a differentiable drift function from $\mathbb{R}^{m}$ into itself satisfying, for some constants $a>0$ and $c\geq 0$,
@@
Moreover, $\tau$ stands for some $(m\times m_{1})$-matrix valued continuous function such that, for all $x\in S$, the matrix $\tau(x)\tau(x)^{\top}$ is invertible and the map $x\mapsto\|\tau(x)\|/\|x\|$ vanishes at infinity. In what follows, for any $x\in S$ we let
$\|\tau(x)\|_{F}^{2}:=\mbox{\rm Tr}(\tau(x)\tau(x)^{\top})\leq m_{1}\|\tau(x)\|_{2}^{2}\quad\text{ with }\|\tau(x)\|_{2}^{2}=\lambda_{\max}(\tau(x)\tau(x)^{\top})$
@@ 3.5.2 The overdamped Langevin diffusion model
An important class of diffusion processes arising in statistical physics and applied probability is the overdamped Langevin diffusion flow $X_{t}(y)$, defined as in (3.30) with $(m_{1},\tau(x))=(m,\sigma I)$ for some $\sigma>0$ and the drift function $b(x)=-\gamma\nabla U(x)$ associated with some parameter $\gamma>0$ and smooth potential function $U$ satisfying the following condition
@@
$W(x):=\exp(\delta\|x\|)\Longrightarrow\sup_{0\leq h^{\prime}\leq 1}\|\widetilde{Q}_{h^{\prime}}(W)/W\|<\infty\text{ and }\widetilde{Q}_{h}(W)/W\leq c_{h}/W^{\delta_{h}}.$ (3.34)
@@ 4 Proofs
showing that if $D_{\psi}(\mu_{1},\mu_{2})>0$ then
@@
Let $x,y\in S$ be such that $x\neq y$ and $r>r_{\epsilon}\vee r_{0}$ be fixed in what follows. In addition, below the functions $(\varpi_{\rho},\varphi_{\rho})$ indexed by $\rho>0$ are as defined in (2.3) and (2.4).
@@
$\begin{split}\ell_{\varphi_{\rho}}(P)(x,y)&\leq\ell_{\varpi_{\rho}}(P)(x,y)=\frac{1+\rho P(V)(x)+\rho P(V)(y)}{1+\rho\varpi_{V}(x,y)}\\ &\leq\frac{1+\rho\big(1+\epsilon\varpi_{V}(x,y)\big)}{1+\rho\varpi_{V}(x,y)}\\ &\leq\epsilon+\frac{1+\rho-\epsilon}{1+\rho r}=1-(1-\epsilon)~\frac{\rho r}{1+\rho r}~\left(1-\frac{r_{\epsilon}}{r}\right)\end{split}$ (4.12)
where the second inequality uses the assumption that (H1) holds with $c=1/2$. Simple calculations show that